# ANIMALBOOTH: MULTIMODAL FEATURE ENHANCEMENT FOR ANIMAL SUBJECT PERSONALIZATION

### v1

@@ 3.2 Qualitative Results
As shown in Fig. 2, methods like BLIP-Diffusion [26], Omnigen [27] and IP-Adapter [15] struggle with maintaining fidelity to the original animal’s identity. For instance, in the ”cheetah” example, Blip introduces noticeable distortions in the cheetah’s facial features and fur patterns. Omnigen, while better, still presents a cheetah that looks somewhat less dynamic and natural compared to the reference. IP-Adapter’s cheetah appears to have some color inaccuracies and a less refined texture. Similar issues are observable in the other examples. For the ”reindeer,” these methods often fail to capture the subtle nuances of its fur, antler structure, or the serene expression seen in the reference, sometimes introducing unnatural postures or color shifts. The ”Highland cow” examples highlight a lack of accurate texture and the distinct shagginess that characterizes the breed. Finally, the ”zebra” images from these comparative methods often fall short in reproducing the sharp stripe patterns, the reflective quality of its fur under moonlight, or the natural stance, sometimes resulting in blurry details or less vibrant contrasts. AnimalBooth consistently demonstrates superior performance in preserving both coarse structures and fine-grained textures. For the cheetah, AnimalBooth accurately reproduces the graceful pose, the distinct spot patterns, and the texture of its fur, capturing the animal’s dynamism. For the reindeer, it faithfully renders the intricate details of its antlers, the texture of its winter coat, and the overall serene atmosphere of the snowy landscape. The Highland cow generated by AnimalBooth accurately showcases the characteristic long, shaggy fur and robust build, maintaining the authenticity of the breed. In the zebra example, AnimalBooth excels at restoring the crispness of the stripes, the luminous sheen of its coat under the moonlight, and a natural, lifelike posture.
3.3 Quantitative Results
As presented in Table 1, AnimalBooth obtained a CLIP-T [22] score of 20.73, which is notably higher than the next best method, Blip [26], at 19.68. This indicates AnimalBooth’s superior ability to comprehend and capture textual semantics, generating animal images that are highly consistent with their descriptions. Concurrently, AnimalBooth achieved the highest DINO [28] score of 75.66, compared to IP-Adapter’s [15] 72.88. This highlights AnimalBooth’s exceptional capability in capturing image details and realism, effectively preserving the intricate textures and fine structures of animals. Furthermore, AnimalBooth also achieved leading scores of 90.00 and 49.08 for CLIP-I and LPIPS, respectively, comprehensively outperforming other methods. This further validates the effectiveness of our proposed Animal-Net and frequency-controlled module in precisely extracting animal identity features and effectively controlling the visual attributes of generated images.
Blip [26] 68.21 62.96 19.68 82.38
Omnigen [27] 71.68 50.05 19.41 72.95
IP-Adapter [15] 62.91 72.88 19.39 89.75
AnimalBooth 49.08 75.66 20.73 90.00
Table 1: Quantitative comparison of different methods.
@@ 3.3 Quantitative Results
Table 2: Ablation study on frequency configurations.
@@ 3.4 Ablation Study
Frequency Configuration. We evaluating the impact of five distinct frequency configuration modes on generated image quality: Low-Pass, Mini-Pass, Mid-Pass, High-Pass, and All-Pass. As presented in Table 2, the Low-Pass frequency configuration significantly outperforms other configurations on CLIP-I, DINO and LPIPS metrics, indicating its superior performance in preserving animal identity details such as fur textures, patterns, and coat colors. This effectively mitigates the detail distortion caused by cross-domain feature space misalignment. In contrast, while High-Pass shows a slight advantage in semantic structure alignment (CLIP-T), its DINO and CLIP-I scores are notably lower, which is detrimental to the fine-grained restoration of individual animal characteristics. Therefore, we prioritize the Low-Pass DCT filtering configuration in AnimalBooth.
Hyper-parameter $\lambda$. In Figure 4 demonstrates the effects of the hyper-parameter $\lambda$ on generated samples with a fixed random seed. As $\lambda$ increases to 1.0, the generated animal gradually loses structural integrity, appearing more diffuse and less defined. A smaller $\lambda$ ensures the generated results adhere more closely to the input animal’s identity, while a larger $\lambda$ biases the results towards the generic diffusion prior. This indicates that $\lambda$ effectively balances precise identity preservation with the model’s broader generative capabilities. Consequently, we empirically set $\lambda$ to 0.4 in our experiments.
Fig. 5: Output results of different frequency filter masks.
3.5 Visulization
Fig. 5 demonstrates the impact of different frequency filter masks on input images, revealing how specific frequency ranges influence the visual outcome. The ”Input” column presents the original high-resolution images. The ”Mini-pass” filter significantly blurs the image by retaining only the extremely low-frequency components, resulting in an abstract visual representation. The ”Low-pass” filter also blurs but less severely, preserving general shapes and smooth color transitions by allowing more low-frequency information. ”Mid-pass” filters isolate intermediate frequencies, accentuating broader textural details and prominent edges. In contrast, ”High-pass” filters emphasize the highest frequencies, sharpening fine details and highlighting sharp edges and intricate textures within the image.
@@ 4 CONCLUSION
This paper introduces AnimalBooth, a fine-tuning-free personalized generation framework specifically designed for animal subjects. Experiments were conducted on our self-constructed AnimalBench dataset, comprising 10,958 training images and 1,000 test images. By integrating the Animal-Net and the adaptive attention module, AnimalBooth achieves state-of-the-art CLIP-T performance. Furthermore, through the low-pass configuration of the DCT frequency-controlled feature integration module, we have enhanced the LPIPS metric to 20 percentage points and effectively improved DINO. Future work will explore extending this framework to animal consistency generation tasks in video modalities.

### v2

@@ 3.2 Qualitative Results
As shown in Fig. 1, methods like BLIP-Diffusion (Li et al., 2023a), Omnigen (Xiao et al., 2025) and IP-Adapter (Ye et al., 2023) struggle with maintaining fidelity to the original animal’s identity. For instance, in the “cheetah” example, BLIP-Diffusion introduces noticeable distortions in the cheetah’s facial features and fur patterns. Omnigen, while better, still presents a cheetah that looks somewhat less dynamic and natural compared to the reference. IP-Adapter’s cheetah appears to have some color inaccuracies and a less refined texture. Similar issues are observable in the other examples. For the “reindeer,” these methods often fail to capture the subtle nuances of its fur, antler structure, or the serene expression seen in the reference, sometimes introducing unnatural postures or color shifts. The “Highland cow” examples highlight a lack of accurate texture and the distinct shagginess that characterizes the breed. Finally, the “zebra” images from these comparative methods often fall short in reproducing the sharp stripe patterns, the reflective quality of its fur under moonlight, or the natural stance, sometimes resulting in blurry details or less vibrant contrasts. AnimalBooth consistently demonstrates superior performance in preserving both coarse structures and fine-grained textures.
Table 1: Ablation on projection layer.
Projection LPIPS $\downarrow$ DINO $\uparrow$ CLIP-T $\uparrow$ CLIP-I $\uparrow$
Concatenation 57.41 31.05 19.85 70.58
MLP Adapter 56.81 31.68 20.02 71.00
Q-Former (Ours) 49.08 75.66 20.73 90.00
Table 2: Ablation on guidance scale $w$.
$w$ LPIPS $\downarrow$ DINO $\uparrow$ CLIP-T $\uparrow$ CLIP-I $\uparrow$
2.0 54.03 43.95 18.92 76.52
5.0 51.26 68.34 20.15 85.67
7.5 (Ours) 49.08 75.66 20.73 90.00
10.0 50.12 72.45 20.58 88.34
Table 3: Efficiency comparison on NVIDIA A100.
Method Time (s) $\downarrow$ VRAM (GB) $\downarrow$ Params (B)
Omnigen (Xiao et al., 2025) 140.0 10.4 3.8
Flux (DiT) (Chang et al., 2024) 85.0 12.8 12.0
AnimalBooth 5.5 0.7 1.2
Table 4: Ablation study on frequency configurations.
@@
3.3 Quantitative Results
As presented in Table 5, AnimalBooth obtained a CLIP-T (Radford et al., 2021) score of 20.73, which is notably higher than the next best method, BLIP-Diffusion (Li et al., 2023a), at 19.68. This indicates AnimalBooth’s superior ability to comprehend and capture textual semantics, generating animal images that are highly consistent with their descriptions. Concurrently, AnimalBooth achieved the highest DINO (Caron et al., 2021) score of 75.66, compared to IP-Adapter’s (Ye et al., 2023) 72.88. This highlights AnimalBooth’s exceptional capability in capturing image details and realism, effectively preserving the intricate textures and fine structures of animals. Furthermore, AnimalBooth also achieved leading scores of 90.00 and 49.08 for CLIP-I and LPIPS, respectively, comprehensively outperforming other methods.
Table 5: Quantitative comparison with state-of-the-art methods.
Textual Inv. (Gal et al., 2022) 72.35 48.92 18.76 71.23
BLIP (Li et al., 2023a) 68.21 62.96 19.68 82.38
Omnigen (Xiao et al., 2025) 71.68 50.05 19.41 72.95
IP-Adapter (Ye et al., 2023) 62.91 72.88 19.39 89.75
Flux (DiT) (Chang et al., 2024) 65.47 55.31 19.52 78.64
AnimalBooth 49.08 75.66 20.73 90.00
Comparison with Optimization-based and DiT Methods. Table 5 also compares AnimalBooth with Textual Inversion (Gal et al., 2022) (a representative optimization-based method) and Flux (Chang et al., 2024) (a state-of-the-art DiT-based model). Textual Inversion, despite requiring per-subject optimization, achieves suboptimal results due to limited embedding capacity for complex animal features. Flux, while generating high-quality images, suffers from identity drift when personalizing animal subjects without specialized adaptation. AnimalBooth outperforms both methods across all metrics while requiring no per-subject optimization at inference time.
Figure 5: Output results of different frequency filter masks.
@@ 3.4 Ablation Study
Frequency Configuration. We evaluate the impact of five distinct frequency configuration modes on generated image quality: Low-Pass, Mini-Pass, Mid-Pass, High-Pass, and without frequency conditioning. As presented in Table 4, the Low-Pass frequency configuration significantly outperforms other configurations on CLIP-I, DINO and LPIPS metrics, indicating its superior performance in preserving animal identity details such as fur textures, patterns, and coat colors. This is because low-frequency components capture the global structure and morphology essential for identity recognition, while high-frequency details can be effectively reconstructed by the diffusion model. In contrast, while High-Pass shows a slight advantage in semantic structure alignment (CLIP-T), its DINO and CLIP-I scores are notably lower, which is detrimental to the fine-grained restoration of individual animal characteristics.
Projection Layer Design. Table 2 compares different projection layer designs. Q-Former significantly outperforms alternatives, improving CLIP-I by 19.42 points and DINO by 44.61 points. This demonstrates Q-Former’s effectiveness as a semantic bottleneck for identity-relevant features.
Guidance Scale. Table 2 presents the ablation on guidance scale $w$. Low guidance ($w=2.0$) produces poor identity preservation. The optimal $w=7.5$ achieves the best balance between identity fidelity and generation quality.
Hyper-parameter $\lambda$. Figure 4 demonstrates the effects of the hyper-parameter $\lambda$ on generated samples with a fixed random seed. As $\lambda$ increases to 1.0, the generated animal gradually loses structural integrity. A smaller $\lambda$ ensures the generated results adhere more closely to the input animal’s identity. Consequently, we empirically set $\lambda$ to 0.4 in our experiments.
Visualization. Fig. 5 demonstrates the impact of different frequency filter masks. The “Mini-pass” filter significantly blurs the image, whereas “Low-pass” preserves general shapes and smooth color transitions, which is critical for identity preservation. “Mid-pass” and “High-pass” filters accentuate textural details and sharp edges respectively.
3.5 Limitations
Despite these advantages, our method has limitations. The reliance on low-frequency structural guidance can occasionally constrain the diversity of generated poses, particularly for highly dynamic actions that differ significantly from the reference. Additionally, while the Q-Former effectively filters background noise, it may occasionally overlook extremely subtle identity cues that are not semantically salient. Future work will explore extending this framework to animal consistency generation tasks in video modalities.
@@ 4 Conclusion
[…]